# The lengths of conjugators in the model filiform groups

### v1

@@ Abstract.
The conjugator length function of a finitely generated group $\Gamma$ gives the optimal upper bound on the length of a shortest conjugator for any pair of conjugate elements in the ball of radius $n$ in the Cayley graph of $\Gamma$. We prove that polynomials of arbitrary degree arise as conjugator length functions of finitely presented groups. To establish this, we analyse the geometry of conjugation in the discrete model filiform groups $\Gamma_{d}=\mathbb{Z}^{d}\rtimes_{\phi}\mathbb{Z}$ where is $\phi$ is the automorphism of $\mathbb{Z}^{d}$ that fixes the last element of a basis $a_{1},\dots,a_{d}$ and sends $a_{i}$ to $a_{i}a_{i+1}$ for $i<d$. The conjugator length function of $\Gamma_{d}$ is polynomial of degree $d$. 2020 Mathematics Subject Classification: 20F65, 20F10, 20F18
@@
The first author thanks the Mathematics Department of Stanford University for its hospitality. The second author gratefully acknowledges the financial support of the National Science Foundation (NSF OIA-2428489).)
@@ 1. Introduction
Whereas Dehn (isoperimetric) functions are the most natural measure of the complexity of a direct approach to the word problem in a finitely presented group, conjugator length functions are the most obvious measure of the complexity of a direct approach to the conjugacy problem in a finitely generated group. By definition, ${\rm{CL}}_{G}(n)$ is the least integer $N$ such that whenever a pair of words $u$ and $v$ on the generators represent conjugate elements of $G$ and the sum of their lengths is at most $n$, there is a word $w$ of length at most $N$ such that $uw=wv$ in $G$. The study of conjugator length functions is far less developed than the study of Dehn functions. In particular, whereas it is essentially known which functions arise as Dehn functions [BORS02], the class of functions that arise as conjugator length functions is poorly understood. Indeed, until recently, very few sharp estimates on conjugator length functions had been established. In part, this reflects the greater delicacy of the conjugacy problem: for example, there exist pairs of finitely generated groups $H<G$ with $|G/H|=2$ such that $H$ has a solvable conjugacy problem and $G$ does not [CM77]. Such examples warn us that, in contrast to the study of Dehn functions, we cannot hope to understand conjugacy length functions using the techniques of coarse geometry. On the other hand, the conjugacy problem is fundamentally geometric in nature: understanding conjugacy in a group $\Gamma$ is closely tied to an understanding of the geometry of annuli with prescribed boundaries in any space with fundamental group $\Gamma$. We refer to our article with Andrew Sale [BRSb] for a more detailed introduction to these ideas and a survey of what is known about conjugator length functions.
In this article we will advance the state of the art by showing that polynomials of arbitrary degree arise as conjugator length functions of finitely presented groups. We shall do so by investigating the geometry of the conjugacy problem in the discrete model filiform groups $\Gamma_{d}$. By definition, $\Gamma_{d}=\mathbb{Z}^{d}\rtimes_{\phi}\mathbb{Z}$, where, with respect to a fixed basis $a_{1},\dots,a_{d}$ for $\mathbb{Z}^{d}$, the automorphism $\phi$ fixes $a_{d}$ and maps $a_{i}\mapsto a_{i}a_{i+1}$ for $i=1,\dots,d-1$.
@@
The corresponding nilpotent Lie groups $G_{d}=\mathbb{R}^{d}\rtimes_{\phi}\mathbb{R}$ have been studied extensively in the context of Carnot geometry, while the lattices $\Gamma_{d}<G_{d}$ have served as key examples for nilpotent and polynomial phenomena in geometric group theory. In particular, $(\Gamma_{d})_{d\in\mathbb{N}}$ was among the first families of groups used to prove Dehn functions can be polynomial of arbitrary degree; see [BMS93, BP94]. Here, we prove that the groups $(\Gamma_{d})_{d\in\mathbb{N}}$ illustrate the same diversity of behaviour among conjugator length functions. In a companion article [BR], we will show that 2-step nilpotent groups can also have conjugator length functions of arbitrary polynomial degree.
@@
Theorem 1.1 provides a counterpoint to our work with Andrew Sale on the conjugator length functions of free-by-cyclic groups: in [BRSa] we proved that ${\rm{CL}}(n)\simeq n$ for the group $F_{d}\rtimes\mathbb{Z}$ obtained by removing the relations $[a_{i},a_{j}]=1$ from the presentation of $\Gamma_{d}$. It also contrasts with Sale’s result [Sal16, Theorem 4.1] that if $\theta$ is diagonalisable over $\mathbb{R}$ and all its eigenvalues have absolute value greater than 1, then $\mathbb{Z}^{d}\rtimes_{\theta}\mathbb{Z}$ has ${\rm{CL}}(n)\simeq n$.
@@
An outline of the proof. Whenever one tries to understand a conjugacy problem in a group $\Gamma$, one is inevitably led to study the structure of centralisers in $\Gamma$, because the set of solutions $x$ to an equation $x^{-1}\gamma_{1}x=\gamma_{2}$ is a coset of $C(\gamma_{1})$. Thus a feature of our study is that we will need a close understanding of the structure of centralisers in $\Gamma_{d}$; this is described in Section 4. It will emerge from this analysis of centralisers that we also need tight control on the geometry of roots in $\Gamma_{d}$, by which we mean solutions to equations $x^{p}=\gamma$; this is the subject of Section 5. With these tools in hand, we will prove Theorem 1.1 by means of an induction on $d$ that exploits the observation that the centre of $\Gamma_{d}$ is infinite cyclic, generated by $a_{d}$, and the quotient of $\Gamma_{d}$ by its centre is isomorphic to $\Gamma_{d-1}$, with the images of $t$ and $a_{i}\ (i<d)$ satisfying the defining the relations of $\Gamma_{d-1}$. When working with this inductive structure, it is important to keep in mind that if two words $u,v$ in the letters $t,a_{1},\dots,a_{d-1}$ define the same element in $\Gamma_{d-1}$, they will in general not be equal in $\Gamma_{d}$; rather, an equality of the form $u=va_{d}^{r}$ will hold in $\Gamma_{d}$. The first lesson to be taken from this observation is that we must be careful to specify in which group equalities between words are taking place. The second point to absorb is that we will have reason to control $|r|$ in expressions of the form $u=va_{d}^{r}$.
@@
With these preparatory thoughts in mind, we can outline our strategy of proof. The actual proof requires us to keep track of various constants, but for the purposes of this outline it makes sense to absord these into symbols such as $\preceq$.
@@
The lower bound $n^{d}\preceq{\rm{CL}}_{\Gamma_{d}}(n)$ is established by arguing that $t^{r}$ is the unique shortest word conjugating $a_{d-1}$ to $a_{d-1}a_{d}^{r}$ in $\Gamma_{d}$ and that $d(1,\,a_{d-1}a_{d}^{n^{d}})\simeq n$ (Section 3). The argument to establish that ${\rm{CL}}_{\Gamma_{d}}(n)\preceq n^{d}$ is more involved. We must show that if $u$ and $g$ are conjugate in $\Gamma_{d}$ and both lie in the ball of radius $n$ about $1$, then there exists $x$ such that $x^{-1}ux=g$ and $|x|\preceq n^{d}$. We will see that the difficult case to deal with is when $g$ does not lie in the normal subgroup $\mathbb{Z}^{d}\rtimes 1$, so let us assume here that this is the case. Assuming that the theorem has been proved for $\Gamma_{d-1}$, we get a word $\theta$ with length $|\theta|\preceq n^{d-1}$ in the letters $t,a_{1},\dots,a_{d-1}$ conjugating the images of $u$ and $g$ in $\Gamma_{d-1}$. This tells us that $\theta^{-1}u\theta=ga_{d}^{r}$ in $\Gamma_{d}$ for some integer $r$. We will use our knowledge of the Dehn function in $\Gamma_{d-1}$ to ensure that $|r|\preceq n^{d(d-1)}$. To complete the induction, we must argue that if $ga_{d}^{r}$ is conjugate to $g$ and $|r|\preceq n^{d(d-1)}$, then the conjugacy can be achieved by a conjugator $h$ with $d(1,h)\preceq n^{d}$. (Note that $ga_{d}^{r}$ may lie outside the ball of radius $n$ in $\Gamma_{d}$, which complicates the structure of the argument.)
@@
It transpires that it is best to realise the conjugacy of $g$ to $ga_{d}^{r}$ in two steps. In the first step, we conjugate by a suitable power of $a_{d-1}$: the aim here is to reduce $|r|$ to something of significantly smaller order by conjugating by $a_{d-1}^{M}$ with $M\approx n^{d(d-1)}$, and since the distortion of $\langle a_{d-1}\rangle<\Gamma_{d}$ is polynomial of degree $d-1$, the length of such a conjugator is roughly $n^{d}$, which is what we are aiming for. We will see that, under the constraint $M\approx n^{d(d-1)}$, we can reduce $|r|$ to something less than $pqe<ne$ where $p$ is the maximal power such that the image of $g$ in $\Gamma_{d-1}$ can be written in the form ${\overline{g}}=g_{0}^{p}a_{d-1}^{*}$ and $q$ is the exponent sum of $t$ in $g_{0}$, and $e=\gcd(p,q)$. A surprisingly large amount of work is required for the final step in the proof, wherein we must bound the length of a conjugator taking $g$ to $ga_{d}^{r}$ for small values of $r$: this involves understanding the action on the coset $g\langle a_{d}\rangle$ of the centraliser of ${\overline{g}}\in\Gamma_{d-1}$ through a homomorphism $\zeta_{{\overline{g}}}:C({\overline{g}})\to\mathbb{Z}$ defined in Definition 7.3. Our understanding of this action relies on an estimate $d(1,g_{0})\preceq d(1,g)$ that comes from the analysis of the lengths of roots in $\Gamma_{d-1}$ in Section 5.
@@ 2. Preliminaries
Throughout, $\Gamma_{d}$ will denote the group $\mathbb{Z}^{d}\rtimes_{\phi}\mathbb{Z}$ where, with respect to a fixed basis, $\phi$ and $\phi^{-1}$ are represented by the following matrices, respectively:
@@
We work with the standard presentation
@@
All distances in $\Gamma_{d}$ will be taken with respect to the generating set of this presentation and the area of any null-homotopic word will be calculated with respect to this set of defining relations. It is worth noting, however, that $\Gamma_{d}$ is generated by just two elements, $a_{d}$ and $t$.
We are interested in the conjugator length function ${\rm{CL}}:\mathbb{N}\to\mathbb{N}$ of these groups. By definition, ${\rm{CL}}(n)$ is the minimal integer $N$ such that whenever a pair of words $u$ and $v$ of length at most $n$ in the generators represent conjugate elements in $\Gamma_{d}$, there is a word $w$ of length at most $N$ such that $uw=wv$ in $\Gamma_{d}$. If we worked with a different finite generating set, the resulting conjugator length function would be equivalent in the sense of the following standard definition.
@@
in the free group $F=F(t,a_{1},\dots,a_{{d-1}})$ with each $\rho_{i}^{\pm 1}$ one of the defining relations of $\Gamma_{d-1}$. We modify the righthand side of this expression by replacing each instance of $[a_{d-1},t]$ with $[a_{d-1},t]a_{d}$, leaving the other relations untouched. Let $\rho_{i}^{\prime}$ be the resulting relations. Then, in $F\times\langle a_{d}\rangle$ we have the equality
@@
where $-r$ is the number of letters $a_{d}$ added during the editing of the $\rho_{i}$. The righthand side has image $a_{d}^{r}$ under the surjection $F\times\langle a_{d}\rangle\to\Gamma_{d}$ that is implicit in the notation. Thus $u=va^{r}$ in $\Gamma_{d}$, with $|r|\leq A$. ∎
@@ 3. Normal forms and distance
The only non-trivial assertion is the one about the size of the exponents. We proceed by induction on $n$. Given a word $w=b_{1}\cdots b_{n}\beta$ of length $n+1$, we apply the inductive hypothesis to write $w^{\prime}=b_{1}\cdots b_{n}$ in normal form, then shuffle $\beta$ past the letters of this normal form to write $w$ in normal form. If $\beta=a_{j}^{\pm 1}$ then this shuffling leaves $r$ and $p_{i}$ unchanged for $i\neq j$ and alters $p_{j}$ by ±$\pm 1$. If $\beta=t$ then this shuffling leaves $p_{1}$ unchanged and for $i\geq 1$ increases $p_{i}$ by $p_{i-1}$, so the observation $n^{i}+n^{i-1}<(n+1)^{i}=n^{i}+in^{i-1}+\cdots$ finishes the induction. If $\beta=t^{-1}$ then the shuffling replaces $p_{i}$ by
@@
Lemma 3.1 and Proposition 3.2 provide the complementary upper and lower bounds for the following statement.
Corollary 3.4.
Remark 3.5.
This corollary played an important role in [Bri99] where it was proved with an argument that is less elementary but which is more concise and provides greater geometric intuition. This argument, which is due to Gromov [Gro93], takes place in the Malcev completion of $\Gamma_{d}$, which is the nilpotent Lie group $G_{d}=\mathbb{R}^{d}\rtimes_{\phi}\mathbb{R}$, where $\Gamma_{d}$ is a lattice. The Lie algebra of $G_{d}$, which is graded, has presentation
This Lie algebra is generated under brackets by $\{X_{1},T\}$, so there is a Carnot-Caratheodory metric on $G_{d}$: this is defined by fixing an inner product on the plane $\langle X_{1},T\rangle$ so that $\{X_{1},T\}$ is orthonormal, transporting this plane around $G_{d}$ by left translation, then defining the distance $d_{cc}(x,y)$ between two points $x,y\in G_{d}$ to be the infimum of the lengths of piecewise smooth curves from $x$ to $y$ that are everywhere tangent to this field of 2-planes, with the length of tangent vectors measured using the transported inner product.
@@
The key observation here is that an arbitrary element $t^{r}x\in\Gamma_{d}{=\mathbb{Z}^{d}\rtimes_{\phi}\mathbb{Z}}$ with $x\in\mathbb{Z}^{d}\times 1$ conjugates $a_{d-1}$ to $a_{d-1}a_{d}^{r}$, with $x$ playing no role. It follows that for each positive integer $r>0$, the unique shortest element of $\Gamma_{d}$ conjugating $a_{d-1}$ to $a_{d-1}a_{d}^{r}$ is $t^{r}$.
In Lemma 3.3 we established the existence of a constant $D_{d}$ such that $a^{n^{d}}$ lies in the ball of radius $D_{d}n$ about $1\in\Gamma_{d}$, hence $d(1,a_{d-1}a_{d}^{n^{d}})\leq D_{d}n+1$. As $t^{n^{d}}$ is the unique shortest element conjugating $a_{d-1}$ to $a_{d-1}a_{d}^{n^{d}}$, we deduce

### v2

@@ Abstract.
The conjugator length function of a finitely generated group $\Gamma$ gives the optimal upper bound on the length of a shortest conjugator for any pair of conjugate elements in the ball of radius $n$ in the Cayley graph of $\Gamma$. We prove that polynomials of arbitrary degree arise as conjugator length functions of finitely presented groups. To establish this, we analyse the geometry of conjugation in the discrete model filiform groups $\Gamma_{d}=\mathbb{Z}^{d}\rtimes_{\phi}\mathbb{Z}$ where $\phi$ is the automorphism of $\mathbb{Z}^{d}$ that fixes the last element of a basis $a_{1},\dots,a_{d}$ and sends $a_{i}$ to $a_{i}a_{i+1}$ for $i<d$. The conjugator length function of $\Gamma_{d}$ is polynomial of degree $d$.
2020 Mathematics Subject Classification: 20F65, 20F10, 20F18
@@
The first author thanks the Mathematics Department of Stanford University for its hospitality. The second author gratefully acknowledges the financial support of the National Science Foundation (NSF GCR-2428489). ORCID: 0000-0002-0080-9059 (MRB) and 0009-0004-3699-0322 (TRR)
@@ 1. Introduction
Whereas Dehn (isoperimetric) functions are the most natural measure of the complexity of a direct approach to the word problem in a finitely presented group, conjugator length functions are the most obvious measure of the complexity of a direct approach to the conjugacy problem in a finitely generated group. By definition, ${\rm{CL}}_{G}(n)$ is the least integer $N$ such that whenever a pair of words $u$ and $v$ of length at most $n$ in the generators represent conjugate elements of $G$, there is a word $w$ of length at most $N$ such that $uw=wv$ in $G$. The study of conjugator length functions is far less developed than the study of Dehn functions. In particular, whereas it is essentially known which functions arise as Dehn functions [BORS02], the class of functions that arise as conjugator length functions is poorly understood. Indeed, until recently, very few sharp estimates on conjugator length functions had been established. In part, this reflects the greater delicacy of the conjugacy problem: for example, there exist pairs of finitely generated groups $H<G$ with $|G/H|=2$ such that $H$ has a solvable conjugacy problem and $G$ does not [CM77]. Such examples warn us that, in contrast to the study of Dehn functions, we cannot hope to understand conjugacy length functions using the techniques of coarse geometry. On the other hand, the conjugacy problem is fundamentally geometric in nature: understanding conjugacy in a group $\Gamma$ is closely tied to an understanding of the geometry of annuli with prescribed boundaries in any space with fundamental group $\Gamma$. We refer to our article with Andrew Sale [BRS26] for a more detailed introduction to these ideas and a survey of what is known about conjugator length functions.
In this article we will advance the state of the art by showing that polynomials of arbitrary degree arise as conjugator length functions of finitely presented groups. We shall do so by investigating the geometry of the conjugacy problem in the discrete model filiform groups $\Gamma_{d}$. By definition, $\Gamma_{d}=\mathbb{Z}^{d}\rtimes_{\phi_{d}}\mathbb{Z}$, where, with respect to a fixed basis $a_{1},\dots,a_{d}$ for $\mathbb{Z}^{d}$, the automorphism $\phi_{d}$ fixes $a_{d}$ and maps $a_{i}\mapsto a_{i}a_{i+1}$ for $i=1,\dots,d-1$.
@@
The corresponding nilpotent Lie groups $G_{d}=\mathbb{R}^{d}\rtimes_{\phi}\mathbb{R}$ have been studied extensively in the context of Carnot geometry, while the lattices $\Gamma_{d}<G_{d}$ have served as key examples for nilpotent and polynomial phenomena in geometric group theory. In particular, $(\Gamma_{d})_{d\in\mathbb{N}}$ was among the first families of groups used to prove Dehn functions can be polynomial of arbitrary degree; see [BMS93, BP94]. Here, we prove that the groups $(\Gamma_{d})_{d\in\mathbb{N}}$ illustrate the same diversity of behaviour among conjugator length functions. In a companion article [BR25b], we will show that 2-step nilpotent groups can also have conjugator length functions of arbitrary polynomial degree. A polynomial upper bound on the conjugator length functions of arbitrary finitely generated nilpotent groups was established much earlier [MMNV22].
@@
Theorem 1.1 provides a counterpoint to our work with Andrew Sale on the conjugator length functions of free-by-cyclic groups: in [BRS] we proved that ${\rm{CL}}(n)\simeq n$ for the group $F_{d}\rtimes\mathbb{Z}$ obtained by removing the relations $a_{i},a_{j}]=1$ from the presentation of $\Gamma_{d}$. It also contrasts with Sale’s result [Sal16, Theorem 5.4] that if $\theta$ is diagonalisable over $\mathbb{R}$ with positive eigenvalues, then $\mathbb{Z}^{d}\rtimes_{\theta}\mathbb{Z}$ has ${\rm{CL}}(n)\simeq n$.
@@
An outline of the proof. Whenever one tries to understand a conjugacy problem in a group $\Gamma$, one is inevitably led to study the structure of centralisers in $\Gamma$, because the set of solutions $x$ to an equation $x^{-1}\gamma_{1}x=\gamma_{2}$ is a coset of the centraliser $C(\gamma_{1})$. Thus a feature of our study is that we will need a close understanding of the structure of centralisers in $\Gamma_{d}$; this is described in Section 4. It will emerge from this analysis of centralisers that we also need tight control on the geometry of roots in $\Gamma_{d}$, by which we mean solutions to equations $x^{p}=\gamma$; this is the subject of Section 5. With these tools in hand, we will prove Theorem 1.1 by means of an induction on $d$ that exploits the observation that the centre of $\Gamma_{d}$ is infinite cyclic, generated by $a_{d}$, and the quotient of $\Gamma_{d}$ by its centre is isomorphic to $\Gamma_{d-1}$, with the images of $t$ and $a_{i}\ (i<d)$ satisfying the defining relations of $\Gamma_{d-1}$. When working with this inductive structure, it is important to keep in mind that if two words $u,v$ in the letters $t,a_{1},\dots,a_{d-1}$ define the same element in $\Gamma_{d-1}$, they will in general not be equal in $\Gamma_{d}$; rather, an equality of the form $u=va_{d}^{r}$ will hold in $\Gamma_{d}$. The first lesson to be taken from this observation is that we must be careful to specify in which group equalities between words are taking place. The second point to absorb is that we will have reason to control $|r|$ in expressions of the form $u=va_{d}^{r}$.
@@
With these preparatory thoughts in mind, we can outline our strategy of proof. The actual proof requires us to keep track of various constants, but for the purposes of this outline it makes sense to absorb these into symbols such as $\preceq$.
The lower bound $n^{d}\preceq{\rm{CL}}_{\Gamma_{d}}(n)$ is established (Section 3) by arguing that $t^{r}$ is the unique shortest word conjugating $a_{d-1}$ to $a_{d-1}a_{d}^{r}$ in $\Gamma_{d}$ and that $d(1,\,a_{d-1}a_{d}^{n^{d}})\simeq n$. En route we explain how the Hilbert-Waring Theorem facilitates a short combinatorial treatment of the metric geometry of $\Gamma_{d}$.
@@
The argument to establish that ${\rm{CL}}_{\Gamma_{d}}(n)\preceq n^{d}$ is more involved. We must show that if $u$ and $v$ are conjugate in $\Gamma_{d}$ and both lie in the ball of radius $n$ about $1$, then there exists $x$ such that $x^{-1}ux=v$ and $|x|\preceq n^{d}$. We will see that the difficult case to deal with is when $v$ does not lie in the normal subgroup $\mathbb{Z}^{d}\rtimes 1$, so let us assume here that this is the case. Assuming that the theorem has been proved for $\Gamma_{d-1}$, we get a word $\theta$ with length $|\theta|\preceq n^{d-1}$ in the letters $t,a_{1},\dots,a_{d-1}$ conjugating the images of $u$ and $v$ in $\Gamma_{d-1}$. This tells us that $\theta^{-1}u\theta=va_{d}^{r}$ in $\Gamma_{d}$ for some integer $r$. We will use our knowledge of the Dehn function in $\Gamma_{d-1}$ to ensure that $|r|\preceq n^{d(d-1)}$. To complete the induction, we must argue that if $va_{d}^{r}$ is conjugate to $v$ and $|r|\preceq n^{d(d-1)}$, then the conjugacy can be achieved by a conjugator $h$ with $d(1,h)\preceq n^{d}$. (Note that $va_{d}^{r}$ may lie outside the ball of radius $n$ in $\Gamma_{d}$, which complicates the structure of the argument.)
@@
It transpires that it is best to realise the conjugacy of $v$ to $va_{d}^{r}$ in two steps. In the first step, we conjugate by a suitable power of $a_{d-1}$: the aim here is to reduce $|r|$ to something of significantly smaller order by conjugating by $a_{d-1}^{M}$ with $M\approx n^{d(d-1)}$, and since the distortion of $\langle a_{d-1}\rangle<\Gamma_{d}$ is polynomial of degree d1d-1, the length of such a conjugator is roughly $n^{d}$, which is what we are aiming for. We will see that, under the constraint $M\approx n^{d(d-1)}$, we can reduce $|r|$ to something less than $pqe<ne$ where $p$ is the maximal power such that the image of $v$ in $\Gamma_{d-1}$ is an element that can be written in the form ${\overline{v}}=v_{0}^{p}a_{d-1}^{*}$ and $q$ is the exponent sum of $t$ in $v_{0}$, and $e=\gcd(p,q)$. A surprisingly large amount of work is required for the final step in the proof, wherein we must bound the length of a conjugator taking $v$ to $va_{d}^{r}$ for small values of $r$: this involves understanding the action on the coset $v\langle a_{d}\rangle$ in $\Gamma_{d}$ of the centraliser of ${\overline{v}}\in\Gamma_{d-1}$ through a homomorphism $\zeta_{{\overline{v}}}:C({\overline{v}})\to\mathbb{Z}$ defined in Definition 7.4. Our understanding of this action relies on an estimate $d(1,v_{0})\preceq d(1,v)$ that comes from the analysis in Section 5 of the lengths of roots in $\Gamma_{d-1}$.
Acknowledgement
We thank the referees for their careful reading and helpful comments.
@@ 2. Preliminaries
Throughout, $\Gamma_{d}$ will denote the group $\mathbb{Z}^{d}\rtimes_{\phi_{d}}\mathbb{Z}$ where, with respect to a fixed basis, $\phi_{d}$ and $\phi^{-1}_{d}$ are represented by the following matrices, respectively:
@@
When there is no danger of ambiguity, we write $\phi$ instead of $\phi_{d}$. We work with the standard presentation
@@
All distances in $\Gamma_{d}$ will be taken with respect to the generating set of this presentation and the area of any null-homotopic word will be calculated with respect to this set of defining relations. It is worth noting, however, that $\Gamma_{d}$ is generated by just two elements, $a_{1}$ and $t$.
We are interested in the conjugator length function $\rm{CL}$$\mathbb{N}\to\mathbb{N}$ of these groups. By definition, ${\rm{CL}}(n)$ is the minimal integer $N$ such that whenever a pair of words $u$ and $v$ of length at most $n$ in the generators represent conjugate elements in $\Gamma_{d}$, there is a word $w$ of length at most $N$ such that $uw=wv$ in $\Gamma_{d}$. If we worked with a different finite generating set, the resulting conjugator length function would be equivalent in the sense of the following standard definition.
@@
in the free group $F=F(t,a_{1},\dots,a_{{d-1}})$ with each $\rho_{i}^{\pm 1}$ one of the defining relations of $\Gamma_{d-1}$. We modify the righthand side of this expression by replacing each instance of $a_{d-1},t]$ with $a_{d-1},t]a_{d}$, leaving the other relations untouched. Let $\rho_{i}^{\prime}$ be the resulting relations. Then, in $F\times\langle a_{d}\rangle$ we have the equality
@@
where r-r is the number of letters $a_{d}$ added during the editing of the $\rho_{i}$. The righthand side has image $a_{d}^{r}$ under the surjection $F\times\langle a_{d}\rangle\to\Gamma_{d}$ that is implicit in the notation. Thus $u=va_{d}^{r}$ in $\Gamma_{d}$, with $|r|\leq A$. ∎
@@ 3. Normal forms and distance
The essential result for us concerning navigating $\Gamma_{d}$ metrically is:
Theorem 3.1.
This theorem played an important role in [Bri99] where it was proved with an argument from [Gui70, Gro93]. This argument takes place in the Malcev completion of $\Gamma_{d}$, which is the nilpotent Lie group $G_{d}=\mathbb{R}^{d}\rtimes\mathbb{R}$ where $\Gamma_{d}$ is a lattice. The Lie algebra of $G_{d}$, which is graded, has presentation
This Lie algebra is generated under brackets by $\{X_{1},T\}$, so there is a Carnot-Caratheodory metric on $G_{d}$: this is defined by fixing an inner product on the plane $\langle X_{1},T\rangle$ so that $\{X_{1},T\}$ is orthonormal, transporting this plane around $G_{d}$ by left translation, then defining the distance $d_{cc}(x,y)$ between two points $x,y\in G_{d}$ to be the infimum of the lengths of piecewise smooth curves from $x$ to $y$ that are almost everywhere tangent to this field of 2-planes, with the length of tangent vectors measured using the transported inner product.
Here we give an alternative, more combinatorial proof of Theorem 3.1. This proof invokes Hilbert’s solution to Waring’s problem, but is otherwise elementary. Lemma 3.2 and Proposition 3.3 below will establish the complementary upper and lower bounds for the theorem.
@@
The only non-trivial assertion is the one about the size of the exponents. We proceed by induction on $n$. Given a word $w=b_{1}\cdots b_{n}\beta$ of length $n+1$, we apply the inductive hypothesis to write $w^{\prime}=b_{1}\cdots b_{n}$ in normal form, then shuffle $\beta$ past the letters of this normal form to write $w$ in normal form. If $\beta=a_{j}^{\pm 1}$ then this shuffling leaves $r$ and $p_{i}$ unchanged for $i\neq j$ and alters $p_{j}$ by $\pm 1$. If $\beta=t$ then this shuffling leaves $p_{1}$ unchanged and for $i\geq 2$ increases $p_{i}$ by $p_{i-1}$, so the observation $n^{i}+n^{i-1}<(n+1)^{i}=n^{i}+in^{i-1}+\cdots$ finishes the induction. If $\beta=t^{-1}$ then the shuffling replaces $p_{i}$ by
@@
The key observation here is that an arbitrary element $t^{r}x\in\Gamma_{d}{=\mathbb{Z}^{d}\rtimes_{\phi}\mathbb{Z}}$ with $x\in\mathbb{Z}^{d}\rtimes 1$ conjugates $a_{d-1}$ to $a_{d-1}a_{d}^{r}$, with $x$ playing no role. It follows that for each positive integer $r>0$, the unique shortest element of $\Gamma_{d}$ conjugating $a_{d-1}$ to $a_{d-1}a_{d}^{r}$ is $t^{r}$.
In Lemma 3.4 we established the existence of a constant $D_{d}$ such that $a_{d}^{n^{d}}$ lies in the ball of radius $D_{d}n$ about $1\in\Gamma_{d}$, hence $d(1,a_{d-1}a_{d}^{n^{d}})\leq D_{d}n+1$. As $t^{n^{d}}$ is the unique shortest element conjugating $a_{d-1}$ to $a_{d-1}a_{d}^{n^{d}}$, we deduce